OpenStudy (anonymous):

On the first 4 tests of his math class, Michael got an average score of 88. What does he need on the next test to have an overall average of 90? could you help me?

OpenStudy (johnweldon1993):

So we know that the mean...is the sum of all the test scores...divided by the number of tests... So we know the mean is 88...and that was using 4 tests...so we have \[\large \frac{x}{4} = 88\] so we can solve for 'x' to be 352 Now if we know he needs an overall score of 90....and that is out of 5 test...we have \[\large \frac{352 + x}{5} = 90\] S lets multiply both sides by 5 \[\large 352 + x = 450\] And now subtract 352 from both sides of the equation \[\large x = 98\] So he would need a 98 on his 5th test to have an overall grade of 90

OpenStudy (anonymous):

nicely done :)

OpenStudy (johnweldon1993):

Why thank you @kelliegirl33 :)

OpenStudy (anonymous):

thank you so much! could you help me with a few more? please

OpenStudy (johnweldon1993):

Sure if I am able to :)

OpenStudy (anonymous):

Kevin has taken 5 quizzes and his average score so far is 84. If he gets 100, a perfect score, on the remaining 3 quizzes, what will his new average be? this would be a different equation, right?

OpenStudy (johnweldon1993):

Alright so partly the same thing here... we know he took 5 quizzes...and the average score was 84....that 84 is the mean...so \[\large \frac{x}{5} = 84\] does that make sense so far?

OpenStudy (anonymous):

yes.

OpenStudy (johnweldon1993):

Now we know...he will take 3 more quizzes...so his total amount of quizzes will be 8 And if he were to get 100 on all 3 of those...that will be 300 more points added to his total... so from the \[\large \frac{x}{5} = 84\] We multiply both sides by 5 to solve for 'x' \[\large x = 420\] So we have that 'x' = 420 Now if he adds 300 more points to that...we will have a total score of 720 right? and that is out of the 8 quizzes he took...so we have \[\large \frac{720}{8} = ?\]

OpenStudy (anonymous):

so his average will be a 90?

OpenStudy (johnweldon1993):

\[\large \color \red{\checkmark}\]

OpenStudy (anonymous):

okay i have two more. I'm going to try to do them, but if i get something wrong wil you correct me?

OpenStudy (anonymous):

William has taken 4 quizzes and his average score so far is 85. If he gets 100, a perfect score, on the remaining 2 quizzes, what will his new average be?

OpenStudy (johnweldon1993):

Of course hun :)

OpenStudy (anonymous):

\[x/4=85\] correct?

OpenStudy (johnweldon1993):

Perfect so far :)

OpenStudy (anonymous):

then add the 200 points to the 340 and you get 540. then \[540/6 = 90\]

OpenStudy (johnweldon1993):

GREAT!! See you don't need me :)

OpenStudy (anonymous):

haha math isnt my best subject. :( but the last one is on average. i don't think it would be the same equation but idk.

OpenStudy (anonymous):

In her chemistry class, Emily took 4 quizzes. Her scores were 93, 90, 96, and 81. What was her average score on the quizzes? i would just add them together and divide by 4 right?

OpenStudy (anonymous):

i got 360/4 =90. is this right? lol

OpenStudy (johnweldon1993):

That is absolutely correct :)

OpenStudy (anonymous):

thank you so much! if i could i would hug you right now! lol

OpenStudy (johnweldon1993):

lol *hug* :)